### v1

@@ C.1 Reference architecture
We use the post-challenge access to the full True catalog and define a new training area that spans from the center of the image up to the right edge, representing a radius of the primary beam. To evaluate only the effect of the training region, we need to preserve an identical number of examples. In practice, we adjust the thickness of the training radius to have a similar training sample size after selection than what we had in the default training area (Sect. 3.4). Our alternative training area is then a rectangular region with $-2.80\leq RA\leq 0.00$ degrees and $-30.12\leq Dec\leq-29.88$ degrees as illustrated in Fig. 27, and that contains 34913 selected sources that follow the flux distribution presented in Fig. 28. We observe that the amount of unselected sources at low apparent flux is much higher than for the reference catalog, which is simply an effect of having sources in all the primary beam sensitivity regimes from 1 to 0.
We train our network using the same configuration as our reference model and only change the training sample, which results in the YOLO-CIANNA-alt-train-ref model. The prediction is also made as before, but we had to alter the SDC1 scorer code to modify the training area definition. Using this modified code, we can obtain a score, presented in Table 8, and a list of matching sources to produce the same diagnostics as for our YOLO-CIANNA-ref catalog. While the scores cannot be directly compared due to the different testing areas, the number of detected sources is relatively similar to our YOLO-CIANNA-ref catalog. The purity is even slightly improved by 0.70%. The histogram of the apparent flux of the detected sources is presented in Fig. 31, exhibiting a similar distribution to our YOLO-CIANNA-ref result. The average source score is also slightly higher, resulting from a better source score that is likely the result of being provided with examples from all the primary beam regimes. Looking at individual fields produces results visually similar to our YOLO-CIANNA-ref model. However, looking at the distribution of false positive detections over the full image field from Fig. 29 demonstrates that our alternative training area results in a detector with more homogeneous detection performances. Overall, it demonstrates that this definition of training sample would have been a better representation of the image context by exhibiting better generalization results than our approach with the extra noise-only regions.
Figure 30: Source flux distribution histogram in log scale and using log-bins for the True catalog of the training region and the training catalog area after a single step of bootstrap refinement.
YOLO-CIANNA-alt-train-ref
@@
Figure 31: Histograms of the sources as a function of their apparent flux using logarithmic bins for the underlying True catalog and predicted sources in our modified testing area. The bottom part of each frame represents the purity and completeness of each bin of the histograms. The top frame is the results for the YOLO-CIANNA-alt-train-ref model, and the bottom frame is the YOLO-CIANNA-alt-train-large-bootstrap model.
@@ C.2 Larger architecture and bootstrap training
An interesting side effect of this alternative training area is that the network training is slightly more stable. From this observation, we tried providing a bit of extra expressivity with less regularization by replacing the second last layer of our default architecture with two $1{\times}1$ convolutional layers, the first one with 3072 filters and a dropout of 30% and the second one with 2048 filters, constructing the YOLO-CIANNA-alt-train-large model. Training this new architecture with our alternative training area only modestly improves the score. This improvement can not be considered significant since it is of the same order as our typical multiple-training-identical-setup variability.
@@
However, this network’s extra expressivity and stability make it suitable for testing the effect of a bootstrap approach for refining our selection function. In the following, we expose the main principle of this method, but it will be presented in more detail in the second paper of the series. The main idea is that instead of deciding which objects from the simulation should be considered detectable, we land this decision on the detector itself. This can be done by using a trained network on our hand-made selection function to make a prediction over the training region. The obtained detections can then be matched with the underlying list of real objects in the simulation to construct a new training sample. We stress that this process is never used to inject a network prediction as a new target directly. The properties from the True catalog sources that were successfully detected are used to define the new training sample. This approach requires stable training, as selecting a prediction before network convergence or when the network overtrains can harm the produced selection function. This approach can be repeated several times using the successively trained network detections to refine the selection function used for the next training.
@@
We tested this approach on our reference model with the default SDC1 training area, but it did not improve the result significantly. With the default training area, any change in our hand-made selection function results in significant variations in the best achievable score. This variability comes mainly from the detections in the intermediate primary beam regime for which the network has to interpolate from the discontinuous training area. This effect is corrected with the alternative training sample, and the bootstrap training approach becomes useful. It is especially relevant in this context where the visibility of a source is strongly dependent on the beam regime, which was less the case when having access only to a central training area with less sensitivity variation. In such a context, having the network provide information on which sources can be considered visible is especially useful.
@@
We apply this approach to our larger architecture and only do one bootstrap step. We present the resulting selection function in Fig. 30. Compared to our handmade custom function, it removes targets for almost all apparent flux values. The most plausible explanation is that it removes blended sources that are impossible to detect. Another effect is to smooth the low-cut flux in a larger flux interval. The score of the resulting model is indicated in Table 8, and the predicted sources flux distribution in Fig. 31. While all the diagnostics show results that are very similar to what we obtain with our reference architecture on our alternative training sample, we observe a significative score improvement of more than $3\%$, correlated with a better flux characterization score with $\bar{s}^{flux}=0.680$. More improvement would likely be achievable by searching the optimum architecture and hyper-parameters setup with this new training sample.
@@
Overall, this approach removes the necessity for a well-defined selection function. This could be especially useful in a highly dimensional input space where defining a selection function by hand can be difficult. With this approach, the starting selection function could even be the catalog of a less performant detection method. Still, this approach requires access to all the possible targets regardless of their visibility, so only real detections are kept at each step. While a simulated dataset is the best-case scenario for this approach, it could still be employed on observed datasets by confirming the proposed detections with additional observations.

### v2

@@ C.1 Reference architecture
We exposed in Sect. 3.5 the issues that arise from having access to only a small subpart of the full image to train our detector. This section explores an alternative definition for the training area that is more representative of the image field variations. We use the post-challenge access to the full True catalog and define a new training area that spans from the center of the image up to the right edge, representing a radius of the primary beam. We do not add extra noise-only regions since they should not be required anymore. To evaluate only the effect of the training region, we adjust the area around the selected radius so the number of selected sources remains similar to our reference case (Sect. 3.4). Our alternative training area is then a rectangular region with $-2.80\leq\textrm{RA}\leq 0.00$ degrees and $-30.12\leq\textrm{Dec}\leq-29.88$ degrees as illustrated in Fig. 27, and that contains 34913 selected sources. Figure 28 represents the apparent flux distribution of the selected sources, which is similar to the distribution we obtained with the default training region.
We train our network using the same configuration as our reference model and only change the training sample, which results in the YOLO-CIANNA-alt-train model. The prediction pipeline remained unchanged, but we altered the SDC1 scorer code to modify the definition of the training area. We obtain the score presented in Table 8 and a list of matching sources to produce the same diagnostics we did on our reference catalog. While the scores cannot be directly compared due to the different testing areas, the number of detected sources is relatively similar to our reference catalog. The purity is even slightly improved by 0.70%. The histogram of the apparent flux of the detected sources is presented in Fig. 30, exhibiting a similar distribution to our reference result. The average source score is also slightly higher, thanks to the presence of sources from all the primary beam regimes in the training sample. Looking at individual fields, this new model produces results that are visually similar to our reference model. However, Fig. 29 shows that the distribution of false positive detections over the full image field is more homogeneous. It demonstrates that this alternative training sample allows for better generalization without requiring extra noise regions.
Alt. training region hand-made selection function
Alt. training region bootstrap selection function
Figure 28: Source flux distribution histogram in log scale and using log-bins for the True catalog and the selected sources in our alternative training. The top frame represents the classical selection function, and the bottom frame represents the bootstrap selection function after a single iteration.
Figure 29: Two-dimensional histograms of various kinds of “objects” representing their distribution over the full SDC1 field for the YOLO-CIANNA-alt-train model. The training area is masked when necessary. All densities are binned using the same 200x200 grid. The match and false detections are made using the DIoU-based matching metric, while the missed ones are based on our selection function. The central coordinates of the field are $\textrm{RA}=0$ deg, $\textrm{Dec}=-30$ deg.
YOLO-CIANNA-alt-train
@@
Figure 30: Histograms of the sources as a function of their apparent flux using logarithmic bins for the underlying True catalog and predicted sources in our modified testing area. The bottom part of each frame represents the purity and completeness of each bin of the histograms. Top and bottom frames represents the alt-train and alt-train-large-bootstrap models respectively.
@@ C.2 Larger architecture and bootstrap training
An interesting side effect of this alternative training area is that the network training is more stable. We tried providing a bit of extra expressivity with less regularization by replacing the second last layer of our default architecture with two $1{\times}1$ convolutional layers, the first one with 3072 filters and a dropout of 30% and the second one with 2048 filters, constructing the YOLO-CIANNA-alt-train-large model. Training this new architecture with our alternative training area only modestly improves the score, as presented in Table 8. This improvement can not be considered significant since it is of the same order as our typical multiple-training-identical-setup variability.
@@
The combination of a more representative training sample and a stronger network makes it possible to use a bootstrap approach for refining our selection function. Instead of deciding which objects from the True catalog should be considered detectable, we land this decision on the detector itself. Using a model trained with a hand-made selection function for a few hundred epochs, we can produce a catalog of predicted sources with an objectness score over the training region. This catalog can then be cross-matched with the underlying list of target objects from the simulation to construct a new training sample. Targets that the detector fails to detect can be removed from the training sample, and targets that are detected outside our hand-made selection function can be added back. We stress that this process is never used to directly inject a detector prediction as a new target. The properties from the True catalog sources that were successfully detected are used to define the new training sample. This approach requires stable training, as selecting a prediction before network convergence or when the network overtrains can harm the self-constrained selection function. This approach can be repeated several times using the successively trained detectors to refine the selection function for the next training.
@@
We tested this approach with our reference architecture using the default SDC1 training area, but it did not improve the result significantly. With the default training area, any change in our hand-made selection function results in significant variations in the best achievable score. This variability mainly comes from the intermediate primary beam regime in which the detector has to interpolate from the discontinuous training area. This effect is corrected with the alternative training area, and the bootstrap training approach actually becomes useful. It is especially relevant in this context where the visibility of a source is strongly dependent on the beam regime, which was less the case for the central training area with less sensitivity variation. In such a context, having the detector provide information on which sources should be detectable is especially useful.
@@
Using our alternative training area and our larger architecture alternative, the bootstrap approach significantly improves the results after only a single step. We present the resulting selection function over the apparent flux histogram in Fig. 28. Compared to our handmade custom function, it removes targets for almost all apparent flux values. The most plausible explanation is that it removes blended sources that the detector cannot separate. Another effect is to smooth the low-cut flux in a wider flux interval. The score of the resulting model is indicated in Table 8, and the predicted sources flux distribution in Fig. 30. While all the diagnostics show results that are very similar to what we obtain with our reference architecture on our alternative training sample, we observe a significative score improvement of more than $3\%$, correlated with a better flux characterization score with $\bar{s}^{\textrm{flux}}=0.680$. More improvement would likely be achievable by searching the optimum architecture and hyperparameters setup with this new training sample.
@@
Overall, this approach removes the necessity for a well-defined selection function. This could be useful in a highly dimensional input space where defining a selection function can be difficult. With this approach, the starting selection function could be the catalog of a less performant detection method. Still, it requires access to all the possible targets regardless of their visibility, so only real detections are kept at each step. While a simulated dataset is a best-case scenario for this approach, it could still be employed on observed datasets by confirming the proposed new detections with additional observations.
Appendix D Impact of simultaneous detection and characterization on detection-only perfomances
In Sect. 5.1.4, we introduce the idea that merging the detection and characterization inside a single loss could result in suboptimal performances. In practice, it is possible to separate both tasks by having a pure object detector and a dedicated characterization network. We could even train an independent characterization network for each parameter to predict. However, we observed that predicting correlated parameters in a single network produces better results for both parameters. For this reason, we expect a combined detection and characterization network to achieve better results as long as the parameter to predict correlates with the detectability of the object. The aim of this section is to verify this point for our method over the SDC1.
To compare models on a detection-only task, we can use the purely geometrical metric defined in Sect. 5.1.4 based on DIoU matching criteria. To evaluate the results at different detection quality requirements we use multiple DIoU threshold values of 0.1, 0.5, and 0.9. To reduce the matching time, predictions below a too-small objectness threshold will be removed. This threshold depends on the detection units and is set to 0.1 for the six small-size regimes and 0.05 for the three intermediate and large-size regimes. This will result in lower AP values than in Sect. 5.1.4, but the respective performances of different models will remain comparable. The results are compiled in Table 9.
One issue with a purely geometrical match is that it increases the chances of a random match compared to the SDC1 match criterium that accounts for the flux. Before training any model, we evaluate the chances of random matches. For this, we build a density map from the raw $\sim$950000 detections from our reference model. This map is used as a prior for drawing an equivalent number of sources with a random position, size, and objectness score that all follow the distributions of our reference model. The resulting “random 950k” catalog is passed through the scoring pipeline, and the results are reported in the same table. The number of random matches for $\textrm{DIoU}>0.5$ is near $3\%$. However, this is not a good representation of our detection catalog as it comprises many detections overlapping over the same object. As stated in Sect. 4.1, almost 80% of the detections are removed by the NMS process that expects interdependency between the detections. Therefore, we build a second random catalog with only 110000 random sources, corresponding to the typical number of objects from our catalog that are not removed and do not match at the end of the post-process pipeline for a DIoU threshold of 0.1. This time, the obtained AP values are likely good approximations of the uncertainty of the AP values over the detected sources catalogs at their respective thresholds.
All the metric scores are reported for our default model with the prediction of the extra parameters as a reference (With param. reference). Our first test deactivates the optimization of the extra-parameters part of the loss while keeping the same architecture and training process (No param.), which is just a toggle switch in our CIANNA framework. As expected, the geometrical score decreases for all DIoU matching values, and the number of detected sources is also reduced. One alternative explanation would be that the network expressivity is now too high for the simplified task, which would also explain why the optimal epoch is found earlier than for our reference model. To test this hypothesis, we train two new models that slightly alter our reference architecture: one where we increase the dropout rate on the before-last layer (No param. more regular), and a second where we reduce the number of filters in the same layer by about 30% (No param. light). Both models perform lower than the simple “No param” case at most DIoU thresholds. This indicates that predicting the extra parameters does have a direct positive impact. Still, we must verify if this effect is generic or if we mitigated the possible negative impact with our quality conditional fitting that results in the extra-parameter cascading loss (Sect. 2.7). We test this hypothesis by retraining our reference model with the extra-parameter prediction activated but setting $L^{{\rm fIoU}}_{p}=-1$ (With param. no cascade). The results are lower than with the parameter deactivated, so predicting parameters for poorly detected sources has a negative impact. In summary, we conclude that extra-parameters can have a positive impact when predicting correlated quantities but only when the negative contribution from unreliable detection is mitigated in the corresponding loss subpart.
Finally, we perform a few extra tests to verify if the lower performance observed when not predicting the extra parameters was only caused by a bad input normalization or augmentation design. For this, we use a per-input patch normalization instead of our global normalization. We train two models with (With param. patch norm) and without (No param. patch norm) the parameter prediction enabled. As expected, the performance of the model with the parameter prediction is lowered, while the performance of the model without the parameter prediction is improved. However, our reference model, which combines parameter prediction and global normalization, remains better. As a last test, we use two input channels corresponding to the two normalization types and train new models with and without the parameter prediction (With param. dual norm. and No param. dual norm.). The detection performances are improved for both models and even surpass our reference model on the detection-only metric, but again, keeping the parameter prediction is better. Still, from the SDC1 metric perspective, our reference score remains higher than with the dual normalization input. This is likely the result of a lack of expressivity in the upper layers of the network that now have to handle two inputs with a limited number of parameters, negatively impacting the source characterization. These results strongly demonstrate that simultaneously predicting correlated parameters with a dynamic loss is a viable approach for improving detection-only performances.
Still, all these results were obtained for a single architecture with only small variations, a single set of hyperparameters, and one dataset. Empirical verification on a per-application basis is advised. Also, we acknowledge that global normalization is only possible in the context of a constant noise image from a single acquisition on a unique instrument. If the objective were to construct a generic source-detector, we would either have to i) input more context information about the instrumental setup or ii) adopt the per-input-patch normalization that forces the detector to base its detection on morphological contrast between the objects and the noise structure or artifacts. The resulting model would likely lose performances in a specific setup but would be more generic and resilient to new structures in the images.
Table 9: Detection-only score metrics for various approaches with and without the extra-parameter part of the loss.
Approach Epoch $\textrm{AP}_{0.1}$ $\textrm{AP}_{0.5}$ $\textrm{AP}_{0.9}$ $N^{m}_{0.1}$ $N^{m}_{0.5}$ $N^{m}_{0.9}$ Score $\textrm{AP}_{SDC1}$ $N^{m}_{SDC1}$
With param. reference 3000 15.71 13.57 0.70 835789 737056 120467 480439 14.47 679778
With param. no cascade 4200 14.94 13.12 0.69 791421 708597 117005 476943 14.24 676288
With param. patch norm. 3200 15.25 13.05 0.71 811351 711355 117859 461417 14.08 654838
With param. dual norm. 3600 15.96 13.74 0.73 857652 751529 126662 473758 14.35 657853
No param. 2600 15.12 13.05 0.58 801494 709853 108664 N/A N/A N/A
No param. more regul. 3000 14.91 13.06 0.73 789935 707086 119189 N/A N/A N/A
No param. light 3000 14.93 12.89 0.67 791478 699148 114422 N/A N/A N/A
No param. patch norm. 2600 15.62 13.18 0.64 836355 723201 113519 N/A N/A N/A
No param. dual norm. 2600 15.84 13.52 0.62 847780 740967 116617 N/A N/A N/A
Random 950k N/A 0.40 0.01 0.00 137877 26416 633 N/A N/A N/A
Random 110k N/A 0.05 0.00 0.00 17152 3306 75 N/A N/A N/A